OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0

OpenStudy (anonymous):

I get sin x = 0 or cos x = 0 but how do I get the solutions from that?

jimthompson5910 (jim_thompson5910):

do you have a unit circle with you?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok to solve sin(x) = 0 for x, you need to find points on the unit circle where the y coordinate is 0

jimthompson5910 (jim_thompson5910):

this is because on the unit circle, any point is of the form (x,y) where x = cos(theta) y = sin(theta) theta = angle

OpenStudy (anonymous):

That would be pi and 2pi

OpenStudy (anonymous):

For cos that would be pi/2 and 3pi/2

jimthompson5910 (jim_thompson5910):

0 is coterminal to 2pi

jimthompson5910 (jim_thompson5910):

so the solutions to sin(x) = 0 are x = 0 and x = pi

jimthompson5910 (jim_thompson5910):

this is on the interval [0, 2pi)

jimthompson5910 (jim_thompson5910):

and good, the solutions to cos(x) = 0 are x = pi/2 and x = 3pi/2

jimthompson5910 (jim_thompson5910):

so in total, the four solutions are x = 0, x = pi, x = pi/2, x = 3pi/2

OpenStudy (anonymous):

\[\frac{ 1 }{2} (2\sin x \cos x)=0,\sin 2x=0=\sin n \pi \] 2x=n pi \[x=\frac{ n \pi }{ 2 }\] for n=0,1,2,3 calculate x

OpenStudy (anonymous):

Got it, thanks

jimthompson5910 (jim_thompson5910):

yw